OpenStudy (anonymous):

The radioactive substance cesium-137 has a half-life of 30 years. The amount A (t) (in grams) of a sample of cesium-137 remaining after t years is given by the following exponential function. A (t)=381 (1/2) ^t/30 Find the initial amount in the sample and the amount remaining after 80 years. Round your answers to the nearest gram as necessary.

OpenStudy (adamaero):

Ignore the second part (for the beginning of setting it up) What is half of cesium-137? 137/2? So... A (30)=381 (1/2) ^(30)/30 A (30)=381 (1/2) ......